A solution of silver nitrate can be used to determine the number of moles of chloride ions in a solution by titration
WHAT is the indicator used in the reaction?
BRIEFLY describe the principle in which the titration is based
1
Expert's answer
2019-08-14T02:43:53-0400

When silver nitrate is added to the solution which contains chloride ions, the white precipitate of silver chloride can be formed according to the reaction below:

Ag+(aq) + Cl-(aq) = AgCl(s).

The indicator, which can be used for this reaction, is potassium chromate (K2CrO4) dilute solution. After all chloride ions have reacted with silver ions, the remaining part (excess) of silver ions (silver nitrate) can react with chromate ions and red - brown precipitate of silver chromate will be formed accodrind to the reaction below:

Ag+(aq) + CrO42-(aq) = Ag2CrO4(s).

This is also known as a Mohr's method.